# Security Analysis of Ponzi Schemes in Ethereum Smart Contracts

### v1

@@ 2.2 Smart Contract
The concept of smart contracts was first proposed by Nick Szabo in 1995, though it remained merely theoretical at the time. Currently, smart contracts have not yet been legally recognized as formal contracts. A smart contract is defined as an automated agreement that is based on blockchain technology. Smart contracts are pieces of computer code that run on the blockchain and manage and enforce the terms of the contract. Smart contracts can automatically execute and enforce the terms and conditions within a contract without requiring a central authority or human intervention, thereby achieving decentralized, automated, and programmable contract execution. Smart contracts can perform various functions, such as issuing and trading digital currencies, asset management, supply chain management, bill settlement, voting, crowdfunding, and more. This provides more flexible and efficient solutions for diverse financial and non-financial applications.
@@ 2.3 Ponzi Scheme
A Ponzi scheme refers to investment fraud in the financial sector, considered the precursor to pyramid schemes. Many illegal multi-level marketing organizations exploit this method to amass wealth. This fraudulent practice was originally devised by a speculator named Charles Ponzi. Simply put, it involves using funds from new investors to pay returns to existing investors, creating a false illusion of quick, substantial profits to lure further investments.
@@ 3.2.1 Tree-Based Ponzi Scheme
Based on the above analysis, it is clear that investors generate profits by recruiting new members. Therefore, to maximize profits, it is essential to extend the contract’s lifespan as much as possible to attract more investors. As shown in Figure 12, the Etheramid contract operated from April 16 to May 11, 2016, with a lifespan of approximately 27 days. Trading activity peaked around April 21. Analysis of its transaction records reveals that 98 accounts were involved in this contract, with 20 accounts achieving a return on investment greater than 1.
@@
There are two vulnerabilities in the Etheramid contract, both of which are SWC-101, indicating the presence of integer overflow vulnerabilities or risks. Furthermore, the analysis indicates that the first problematic function is fallback. Transactions exploiting this vulnerability consume Gas between 2741 and 37542. The consequence of this flaw is that the return value of a message call is not checked, and external calls return a Boolean value. If the callee halts with an exception, ’false’ is returned, and execution continues in the caller. The caller should check whether an exception happened and react accordingly to avoid unexpected behavior. For example, it is desirable to wrap external calls in require() so the transaction is reverted if the call fails. Mythril specifically highlights the problematic line at msg.sender.send(msg.value) on line 57, indicating that the function storing the investor’s address fails to check the return value, creating a vulnerability.
@@ 3.2.2 Chain-Based Ponzi Scheme
It is based on array structures, inducing a linear storage sequence among users. This scheme also represents a special case of tree-based Ponzi schemes. Consequently, apart from the storage structure, its other characteristics closely resemble tree-based Ponzi schemes. However, such schemes typically predefine a constant multiplier applied to users’ investment amounts, with this multiplier being identical for all participants. This scheme first collects all investments until a predetermined amount is reached, then begins disbursing payments to users, passing the chain structure to the next user.
The CrystalDoubler contract is a classic example of this type of scam, and the following is its core code. User investments trigger the enter() function via msg.value. The contract requires a minimum investment of 500 fenny. If less than 500 fenny, the contract rejects the user. Upon successful entry, the user’s address is stored in the array. After a successful investment, the contract updates its own balance. As new investments continue, the contract constantly checks whether the balance is sufficient to pay the returns to users in the array. If sufficient, the contract pays the current investor double their investment amount. This step loops, paying out to investors awaiting returns until the balance no longer meets the payout condition.
@@ 3.2.3 Waterfall-Based Ponzi Scheme
It employs a user ranking similar to the chain-based Ponzi schemes mentioned above, differing only in the logic of currency profit distribution. A chain structure stores each new investor, forming an investor chain. This ensures every investor receives their share when the contract pays out. However, given that payments are made on a first-come, first-served basis, with each distribution commencing from the head of the chain, there is a possibility that investors positioned towards the tail of the chain may never receive profits.
The PonzICO contract is a classic example of this type of scam, and the following is its core code. When total investments fall below 200,000 ETH, the contract first allocates 50% of incoming funds to the contract owner. The remaining funds are then distributed to prior investors. During this process, if sufficient funds exist to pay the first investor in the chain, that investor receives a proportionate return on their original investment. Finally, the contract continues disbursing funds along the investor chain to the next eligible investor, repeating this process until the remaining balance is insufficient for further payments. There are no obvious vulnerabilities in the contract through Mythril analysis, though there remains the possibility of some deeply hidden or complex vulnerabilities within PonzICO.
@@ 3.2.4 Transfer-Based Ponzi Scheme
This scheme determines investor entry fees via smart contracts, with fees increasing as new participants join. At any given moment, only one investor receives payouts. Upon receiving payouts, the investor must relinquish their permission to the next user, a process termed permission transfer. To join the scheme, a user must send a certain amount of ETH to the contract to trigger the payout function. The previous user receives this payment from the contract, records the new user’s address, and doubles the entry fee.
@@ 3.3 The Fomo3D Ponzi Game
Fomo3D is a Ponzi scheme game on Ethereum, which is considered to be a form of gambling. The game uses KEY, the token in the Fomo3D game, and the last player to purchase KEY can win a massive prize. From a gameplay perspective, and taking into consideration the rewards structure, Fomo3D is transparent in its approach, openly informing players that it functions as a Ponzi scheme, a Ponzi fund. Interestingly, Fomo3D’s domain name is exitscam.me, which breaks down to “exit scam me”, translating to “leave the scam me”. Its smart contract code is open-source, making it a refreshing exception among blockchain gambling games. This approach serves to provide players with the assurance that there are no rigged operators present, nor any clandestine entities that will accept donations and subsequently misappropriate them.
In order to participate in Fomo3D, it is first necessary to transfer a specific quantity of ETH to the designated contract address. Following this, the participant is required to select their preferred team from among the options of Snek, Bull, Whale, or Bear. Subsequently, a specific quantity of KEY will be received at the prevailing exchange rate. It has been determined that 86% of the ETH deposited by participants will be allocated to the prize pool and dividends. However, it should be noted that the distribution ratio varies according to team selection, as illustrated in Figure 13. A proportion of 10% is allocated as a form of remuneration to those who successfully refer a new customer to the organization. In the absence of a referrer, the reward is distributed to all KEY holders. A proportion of 2% is allocated to the game project team, 1% is designated for Pot Swap, and 1% is allocated for random airdrops. It is noteworthy that each purchase exceeding 0.01 ETH increases the probability of a random airdrop by 0.1%.
@@
When someone purchases a pass, a 24-hour countdown to win automatically begins. If no one else buys a pass within that 24-hour period, the last person to purchase a pass becomes the game’s winner and receives 48% of the prize pool. However, if another purchase occurs within those 24 hours, the countdown is reset and a further 30 seconds is added. This winner mechanism invariably compels numerous players to purchase tickets during the final 30 seconds of the countdown, hoping to be the final purchaser. After the advent of Fomo3D, a plethora of analogous games have emerged, all of which require players to purchase game items with ETH. In these games, the triumphant player is entitled to claim the so-called “grand prize”, while regular participants have the opportunity to obtain “airdrop rewards”.
This game is governed by smart contracts, which means it can never truly be stopped. When Fomo3D first launched in 2018, many feared it would snowball out of control, potentially crashing Ethereum. This concern stemmed from the risk that excessive ETH investment by players seeking victory could flood the Fomo3D pool, causing a critical shortage of circulating ETH on the Ethereum network. Fortunately, although Fomo3D once reached nearly 30,000 ETH, flaws in the contract code allowed hackers with advanced attack techniques to increase their chances of receiving “airdrop rewards” and even use special methods to claim the “grand prize”. Due to severe exploitation of these two reward mechanisms by hackers, the game’s appeal to users diminished, failing to attract new capital. This ultimately created a vicious cycle, rendering it incapable of significantly impacting Ethereum.
Of all funds entering the game, specifically ETH, 1% flows into the secondary prize pool. The airdrop probability starts at 0%. Each order exceeding 0.1 ETH increases the airdrop probability by 0.1%, capped at 1%. Players who purchase orders between 0.1 and 1 ETH have a chance to win one-quarter of the secondary prize pool, with higher purchase amounts increasing their probability. The Fomo3D airdrop reward mechanism contained two vulnerabilities. First, the “random number” within the contract was not truly random but predictable. Second, there was a flaw in the logic that determined whether the caller was a contract address.
@@
(2) “Contract Address Verification” Vulnerability: To prevent automated attacks on the contract, Fomo3D developers specifically created the isHuman() function to prohibit players from predicting winning random numbers within the contract. Within the isHuman() function, the developers made an error. The extcodesize operator retrieves the code size at a target address. However, for successfully deployed contracts, their addresses correspond to specific code, causing extcodesize to always return a value greater than zero. Consequently, this method could be exploited to determine whether a target address belonged to a contract. The developers of the Fomo3D game evidently employed a similar approach to prevent contracts from calling specific functions. However, they likely overlooked a glaring vulnerability in this method. By calling the game participation function in the contract construction method, one can bypass the restrictions set by the developer. This vulnerability arises because, during the process of constructing the contract, the contract address does not yet correspond to any code, meaning extcodesize returns 0 at this stage.
@@
By exploiting these two vulnerabilities in combination, attackers can create malicious contracts to participate in the game. Leveraging their ability to predict random numbers, they can significantly boost their win rate and profit with impunity. The Mythril tool did not detect any obvious vulnerabilities in the Fomo3D contract. This demonstrates that Mythril is not infallible and cannot detect all vulnerabilities. Therefore, when using program analysis methods to detect Ponzi scheme smart contracts, it is advisable to supplement tool-based analysis with source code review.
@@ 4.1 Experimental Data Collection
The smart contract source code data utilized for batch detection was obtained through a process of data collection from the Etherscan.io website using web crawlers. Since the last smart contract flagged as a Ponzi scheme on Etherscan.io predates 2019, all crawled smart contract source code originates from before 2019, making this comparison more meaningful. The scraped data contains 500 smart contracts.
@@
When crawling webpages, set the retry mechanism to 25. This primarily prevents data download failures caused by network fluctuations during crawling and improves crawling efficiency. Data parsing refers to the process of extracting target data from formats such as HTML, XML, and JSON. After obtaining the webpage content from the URL, simulate browser header information to send messages to Etherscan.io. Finally, save the contract source code retrieved from Etherscan.io as sol files. Name each file according to the crawl sequence, then store all sol files in the code directory.
@@ 4.2 The Batch Detection Process
Mythril smart contract analysis tools can perform detection and analysis on individual smart contracts as well as multiple smart contracts simultaneously, enabling batch detection. Batch detection in Mythril can be performed in two ways: shell scripts and command mode.

### v2

@@ 2.2 Smart Contract
The concept of smart contracts was first proposed by Nick Szabo in 1995, though it remained merely theoretical at the time. Currently, smart contracts have not yet been legally recognized as formal contracts. A smart contract is defined as an automated agreement that is based on blockchain technology. Smart contracts are pieces of computer code that run on the blockchain and manage and enforce the terms of the contract. Smart contracts can automatically execute and enforce the terms and conditions within a contract without requiring a central authority or human intervention, thereby achieving decentralized, automated, and programmable contract execution [46]. Smart contracts can perform various functions, such as issuing and trading digital currencies, asset management, supply chain management, bill settlement, voting, crowdfunding, and more. This provides more flexible and efficient solutions for diverse financial and non-financial applications.
@@ 2.3 Ponzi Scheme
A Ponzi scheme refers to investment fraud in the financial sector, considered the precursor to pyramid schemes. Many illegal multi-level marketing organizations exploit this method to amass wealth [50]. This fraudulent practice was originally devised by a speculator named Charles Ponzi. Simply put, it involves using funds from new investors to pay returns to existing investors, creating a false illusion of quick, substantial profits to lure further investments.
@@ 3.2.1 Tree-Based Ponzi Scheme
Based on the above analysis, it is clear that investors generate profits by recruiting new members. Therefore, to maximize profits, it is essential to extend the contract’s lifespan as much as possible to attract more investors. As shown in Figure 12, the Etheramid contract operated from April 16 to May 11, 2016, with a lifespan of approximately 27 days. Trading activity peaked around April 21. Analysis of its transaction records reveals that 98 accounts were involved in this contract, with 20 accounts achieving a return on investment greater than 1 [49].
@@
There are two vulnerabilities in the Etheramid contract, both of which are SWC-101, indicating the presence of integer overflow vulnerabilities or risks. Furthermore, the analysis indicates that the first problematic function is fallback. Transactions exploiting this vulnerability consume Gas between 2741 and 37542. The consequence of this flaw is that the return value of a message call is not checked, and external calls return a Boolean value. If the callee halts with an exception, ’false’ is returned, and execution continues in the caller. The caller should check whether an exception happened and react accordingly to avoid unexpected behavior. For example, it is desirable to wrap external calls in require() so the transaction is reverted if the call fails. Mythril specifically highlights the problematic line at msg.sender.send(msg.value) on line 57, indicating that the function storing the investor’s address fails to check the return value, creating a vulnerability [51].
@@ 3.2.2 Chain-Based Ponzi Scheme
It is based on array structures, inducing a linear storage sequence among users. This scheme also represents a special case of tree-based Ponzi schemes [35]. Consequently, apart from the storage structure, its other characteristics closely resemble tree-based Ponzi schemes. However, such schemes typically predefine a constant multiplier applied to users’ investment amounts, with this multiplier being identical for all participants. This scheme first collects all investments until a predetermined amount is reached, then begins disbursing payments to users, passing the chain structure to the next user.
The CrystalDoubler contract is a classic example of this type of scam, and the following is its core code. User investments trigger the enter() function via msg.value. The contract requires a minimum investment of 500 fenny. If less than 500 fenny, the contract rejects the user. Upon successful entry, the user’s address is stored in the array. After a successful investment, the contract updates its own balance [24]. As new investments continue, the contract constantly checks whether the balance is sufficient to pay the returns to users in the array. If sufficient, the contract pays the current investor double their investment amount. This step loops, paying out to investors awaiting returns until the balance no longer meets the payout condition.
@@ 3.2.3 Waterfall-Based Ponzi Scheme
It employs a user ranking similar to the chain-based Ponzi schemes mentioned above, differing only in the logic of currency profit distribution [35]. A chain structure stores each new investor, forming an investor chain. This ensures every investor receives their share when the contract pays out. However, given that payments are made on a first-come, first-served basis, with each distribution commencing from the head of the chain, there is a possibility that investors positioned towards the tail of the chain may never receive profits.
The PonzICO contract is a classic example of this type of scam, and the following is its core code. When total investments fall below 200,000 ETH, the contract first allocates 50% of incoming funds to the contract owner [25]. The remaining funds are then distributed to prior investors. During this process, if sufficient funds exist to pay the first investor in the chain, that investor receives a proportionate return on their original investment. Finally, the contract continues disbursing funds along the investor chain to the next eligible investor, repeating this process until the remaining balance is insufficient for further payments. There are no obvious vulnerabilities in the contract through Mythril analysis, though there remains the possibility of some deeply hidden or complex vulnerabilities within PonzICO.
@@ 3.2.4 Transfer-Based Ponzi Scheme
This scheme determines investor entry fees via smart contracts, with fees increasing as new participants join. At any given moment, only one investor receives payouts. Upon receiving payouts, the investor must relinquish their permission to the next user, a process termed permission transfer. To join the scheme, a user must send a certain amount of ETH to the contract to trigger the payout function. The previous user receives this payment from the contract, records the new user’s address, and doubles the entry fee [34].
@@ 3.3 The Fomo3D Ponzi Game
Fomo3D is a Ponzi scheme game on Ethereum, which is considered to be a form of gambling. The game uses KEY, the token in the Fomo3D game, and the last player to purchase KEY can win a massive prize [27]. From a gameplay perspective, and taking into consideration the rewards structure, Fomo3D is transparent in its approach, openly informing players that it functions as a Ponzi scheme, a Ponzi fund. Interestingly, Fomo3D’s domain name is exitscam.me, which breaks down to “exit scam me”, translating to “leave the scam me” [12]. Its smart contract code is open-source, making it a refreshing exception among blockchain gambling games. This approach serves to provide players with the assurance that there are no rigged operators present, nor any clandestine entities that will accept donations and subsequently misappropriate them.
In order to participate in Fomo3D, it is first necessary to transfer a specific quantity of ETH to the designated contract address. Following this, the participant is required to select their preferred team from among the options of Snek, Bull, Whale, or Bear [55]. Subsequently, a specific quantity of KEY will be received at the prevailing exchange rate. It has been determined that 86% of the ETH deposited by participants will be allocated to the prize pool and dividends. However, it should be noted that the distribution ratio varies according to team selection, as illustrated in Figure 13. A proportion of 10% is allocated as a form of remuneration to those who successfully refer a new customer to the organization. In the absence of a referrer, the reward is distributed to all KEY holders. A proportion of 2% is allocated to the game project team, 1% is designated for Pot Swap, and 1% is allocated for random airdrops. It is noteworthy that each purchase exceeding 0.01 ETH increases the probability of a random airdrop by 0.1%.
@@
When someone purchases a pass, a 24-hour countdown to win automatically begins [31]. If no one else buys a pass within that 24-hour period, the last person to purchase a pass becomes the game’s winner and receives 48% of the prize pool. However, if another purchase occurs within those 24 hours, the countdown is reset and a further 30 seconds is added. This winner mechanism invariably compels numerous players to purchase tickets during the final 30 seconds of the countdown, hoping to be the final purchaser. After the advent of Fomo3D, a plethora of analogous games have emerged, all of which require players to purchase game items with ETH. In these games, the triumphant player is entitled to claim the so-called “grand prize”, while regular participants have the opportunity to obtain “airdrop rewards” [14].
This game is governed by smart contracts, which means it can never truly be stopped. When Fomo3D first launched in 2018, many feared it would snowball out of control, potentially crashing Ethereum. This concern stemmed from the risk that excessive ETH investment by players seeking victory could flood the Fomo3D pool, causing a critical shortage of circulating ETH on the Ethereum network [36]. Fortunately, although Fomo3D once reached nearly 30,000 ETH, flaws in the contract code allowed hackers with advanced attack techniques to increase their chances of receiving “airdrop rewards” and even use special methods to claim the “grand prize”. Due to severe exploitation of these two reward mechanisms by hackers, the game’s appeal to users diminished, failing to attract new capital. This ultimately created a vicious cycle, rendering it incapable of significantly impacting Ethereum.
Of all funds entering the game, specifically ETH, 1% flows into the secondary prize pool. The airdrop probability starts at 0%. Each order exceeding 0.1 ETH increases the airdrop probability by 0.1%, capped at 1% [58]. Players who purchase orders between 0.1 and 1 ETH have a chance to win one-quarter of the secondary prize pool, with higher purchase amounts increasing their probability. The Fomo3D airdrop reward mechanism contained two vulnerabilities. First, the “random number” within the contract was not truly random but predictable. Second, there was a flaw in the logic that determined whether the caller was a contract address.
@@
(2) “Contract Address Verification” Vulnerability: To prevent automated attacks on the contract, Fomo3D developers specifically created the isHuman() function to prohibit players from predicting winning random numbers within the contract. Within the isHuman() function, the developers made an error. The extcodesize operator retrieves the code size at a target address. However, for successfully deployed contracts, their addresses correspond to specific code, causing extcodesize to always return a value greater than zero. Consequently, this method could be exploited to determine whether a target address belonged to a contract [40]. The developers of the Fomo3D game evidently employed a similar approach to prevent contracts from calling specific functions. However, they likely overlooked a glaring vulnerability in this method. By calling the game participation function in the contract construction method, one can bypass the restrictions set by the developer. This vulnerability arises because, during the process of constructing the contract, the contract address does not yet correspond to any code, meaning extcodesize returns 0 at this stage.
@@
By exploiting these two vulnerabilities in combination, attackers can create malicious contracts to participate in the game. Leveraging their ability to predict random numbers, they can significantly boost their win rate and profit with impunity [7]. The Mythril tool did not detect any obvious vulnerabilities in the Fomo3D contract. This demonstrates that Mythril is not infallible and cannot detect all vulnerabilities. Therefore, when using program analysis methods to detect Ponzi scheme smart contracts, it is advisable to supplement tool-based analysis with source code review.
@@ 4.1 Experimental Data Collection
The smart contract source code data utilized for batch detection was obtained through a process of data collection from the Etherscan.io website using web crawlers [6]. Since the last smart contract flagged as a Ponzi scheme on Etherscan.io predates 2019, all crawled smart contract source code originates from before 2019, making this comparison more meaningful. The scraped data contains 500 smart contracts.
@@
When crawling webpages, set the retry mechanism to 25. This primarily prevents data download failures caused by network fluctuations during crawling and improves crawling efficiency. Data parsing refers to the process of extracting target data from formats such as HTML, XML, and JSON [26]. After obtaining the webpage content from the URL, simulate browser header information to send messages to Etherscan.io. Finally, save the contract source code retrieved from Etherscan.io as sol files. Name each file according to the crawl sequence, then store all sol files in the code directory.
@@ 4.2 The Batch Detection Process
Mythril smart contract analysis tools can perform detection and analysis on individual smart contracts as well as multiple smart contracts simultaneously, enabling batch detection. Batch detection in Mythril can be performed in two ways: shell scripts and command mode [54].